For the following function f(x)=3x2-x+4 find
f(2)
F(2x)
F(2x+2)


Answer by stanbon(57424)

For the following function
f(x)=3x^2-x+4 find
f(2)= 3*2^2 - 2 + 4 = 12+2 = 14
---------------------------
F(2x) = 3(2x)^2 - 2x + 4 = 12x^2 -2x +4
---------------------------
F(2x+2) = 3(2x+2)^2 + 2x+2 + 4
= 3(4x^2+8x+4) + 2x^2 + 4
---
= 6x^2 + 24x + 16
====================
Cheers,
Stan H.